June 7, 2016, Analysis Working Group

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

The meeting will be on Tuesday June 7, 2016 at 1:30 pm EDT. For those people at Jefferson Lab, the meeting will be in room F326.

Meeting Connections

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/115815824.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 115815824.
  3. To join via phone, use one of the following numbers and the Conference ID: 115815824.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.


  1. Announcements
  2. REST Production Preparations:
  3. Efficiency / Resolution Studies Running with this Launch
    • Tracking: Paul, Simon
    • BCAL Hadronic: Elton
    • FCAL Hadronic: ???
    • SC: Mahmoud
    • TOF: Beni?
    • CDC: Mike
    • FDC: Alex A.
  4. TPOL polarization update - Mike and Nathan
  5. AOB


  1. Announcements
    • DNP Abstracts reminder. There may be 10-15 speakers possible in the mini-symposium.
  2. REST Production Preparations:
    • Spring 2016 Dataset Summary and Trigger Summary:
    • First priority run block: 11347-11555
      • Many updates to CCDB and sim-recon in preparation for this launch
      • Test jobs submitted early Tuesday morning and were being analyzed at the time of the meeting, with a few minor bugs being fixed
    • Second priority run block: 11556-11678
      • After the meeting Sean found that the RF prescale was unstable for some runs during this block (11556-11
      • Calibrations and detector group status reports
      • TAGH and PS: Nathan has analyzed the PS skims for this run block and is studying time offsets
      • TAGM: Alex reported that during this run block the TAGM used the improve bias voltages so the TAGM integral cuts are much simpler, and initial TDC/ADC time offsets are now in the CCDB
      • CDC: Mike is working on time-to-distance tables for the higher magnetic field, and expects to have them in a few days
      • CDC: Naomi prepared new gain tables, and they are now in the CCDB
      • FCAL: Adesh has new gain constants for this run block and is working on the non-linearity correction that appears to be different from the earlier run block
      • BCAL: Will reported that the BCAL gains are ready for production. Andrei is working the BCAL TDC calibration, which will take a few weeks, so we will not include it in this round of production
      • Goal: all calibrations to CCDB for this run block by 5 pm on Friday 6/10. Allows time for testing before expected launch next week
  3. Efficiency / Resolution Studies Running with this Launch
    • Paul reported on the efficiency plugins which are running during the current launch, and that ROOT TTrees are being produced for these studies
  4. TPOL polarization update - Mike and Nathan
    • Mike presented a shortened version of his talk at Monday's BLTWG meeting
    • The results from Nathan and Mike's independent analyses are consistent in the range of 7.9-9.1 GeV
    • There was some discussion about the comparison to the rho beam asymmetry
      • Matt asked if there was a physical mechanism for fine structure in the polarization vs energy, and Richard confirmed that it could not be a crystal effect. More study is needed to see if the fine structure in the polarization vs energy is statistically significant, and if that implies some remaining systematic uncertainty
      • It was suggested to compare the TPOL asymmetry with and without requiring a coincident hit in the tagger where the tagged asymmetry is likely a better comparison to rho asymmetry, albeit with lower statistics
  5. AOB